# UAlign: Pushing the Limit of Template-free Retrosynthesis Prediction with Unsupervised SMILES Alignment

### v1

@@ Abstract
Retrosynthesis planning poses a formidable challenge in the organic chemical industry, particularly in pharmaceuticals. Single-step retrosynthesis prediction, a crucial step in the planning process, has witnessed a surge in interest in recent years due to advancements in AI for science. Various deep learning-based methods have been proposed for this task in recent years, incorporating diverse levels of additional chemical knowledge dependency. This paper introduces UAlign, a template-free graph-to-sequence pipeline for retrosynthesis prediction. By combining graph neural networks and Transformers, our method can more effectively leverage the inherent graph structure of molecules. Based on the fact that the majority of molecule structures remain unchanged during a chemical reaction, we propose a simple yet effective SMILES alignment technique to facilitate the reuse of unchanged structures for reactant generation. Extensive experiments show that our method substantially outperforms state-of-the-art template-free and semi-template-based approaches. Importantly, Our template-free method achieves effectiveness comparable to, or even surpasses, established powerful template-based methods. Scientific contribution: We present a novel graph-to-sequence template-free retrosynthesis prediction pipeline that overcomes the limitations of Transformer-based methods in molecular representation learning and insufficient utilization of chemical information. We propose an unsupervised learning mechanism for establishing product-atom correspondence with reactant SMILES tokens, achieving even better results than supervised SMILES alignment methods. Extensive experiments demonstrate that UAlign significantly outperforms state-of-the-art template-free methods and rivals or surpasses template-based approaches, with up to 5% (top-5) and 5.4% (top-10) increased accuracy over the strongest baseline.
@@ 1 Introduction
Retrosynthesis prediction is crucial in organic chemistry, aiding in finding efficient synthetic pathways from target molecules to accessible starting materials. Despite significant advancements in chemical synthesis technology, it still remains a challenge in industries like pharmaceuticals. The extensive search space and the incomplete understanding of chemical reaction mechanisms make retrosynthesis prediction difficult, even for experienced chemists. To address this issue, computer-assisted synthetic planning (CASP) has gained increasing attention in recent years, starting from the seminal work by Corey. This paper focuses on single-step retrosynthesis prediction, which is the fundamental step in CASP. It aims to predict the reactants that can lead to a given product molecule through a single reaction step.
Various deep-learning-based single-step retrosynthesis prediction methods have been proposed in recent years. These methods can be broadly classified into three groups based on their dependency on additional chemical knowledge: template-based, semi-template-based and template-free methods. Template-based methods [6, 35, 4, 2] requires an extra database of reaction templates. They frame the retrosynthesis prediction as a classification or retrieval problem for reaction templates suitable for the given product molecule to be predicted. Retrosim [4] utilizes molecular similarity to rank reaction templates. LocalRetro [2] and GLN [6] use graph neural networks to model the relationship between reaction templates and molecules to predict the most suitable reaction template. RetroKNN [35] further improves upon LocalRetro by addressing the issue of data imbalance using K-nearest neighbors (KNN). Template-based methods have strong interpretability and can accurately predict reactants. However, template-based methods are often unable to cover all cases and suffer from poor scalability due to limitations imposed by the template database.
@@
Semi-template-based methods to a certain extent are more in line with chemical intuition. However, these methods increase the complexity of inference and training as they break down retrosynthesis into two subtasks. Errors in synthon prediction directly affect subsequent reactant completion and overall performance. Besides, methods based on leaving group necessitates an extra leaving group database. This requirement, akin to template-based approaches, imposes limitations on the model’s scalability.
@@ 2.2 EGAT+
The initial node features $h_{u}^{(0)}$ and edge features $e_{u,v}^{(0)}$ are determined via several chemical property descriptors, whose details are shown in Appendix F. After $K$ iterations of message passing, we can obtain the encoded features $h^{(K)}_{u}$ of all atoms and make up the output $H\in\mathbb{R}^{V_{P}\times d}$ of the encoder, where $d$ denotes the embedding size.
@@ 2.3 SMILES Alignment
For further discussion, we denote the order-preserving reactant SMILES given the reactant molecules $\mathcal{R}$ and a DFS order $O$ of product as $OPSmiles(\mathcal{R},O)$. An example of the process to generate order-preserving reactants SMILES is shown in Fig. 2. The detailed implementations are presented in Appendix E.1.
@@ 3.1 Evaluation Protocol
Benchmark Datasets. We adopt three datasets for evaluation: (1) USPTO-50K consists of 50,016 atom-mapped reactions grouped into 10 different classes; (2) USPTO-FULL comprises 1,013,118 atom-mapped reactions without any reaction class information. (3) USPTO-MIT consists of 479,035 atom-mapped reactions without any reaction class information. To ensure a fair comparison, we adopt the same training/validation/test splits as those in a previous study [6] for USPTO-50K and USPTO-FULL datasets. The training/validation/test splits is aligned with the previous study [11]. The detailed data processing procedure and the statistical information of the processed dataset are presented in Appendix B.
Metrics. We utilize the following three evaluation metrics for evaluation: top-$k$ accuracy, top-$k$ SMILES validity, and top-$k$ round-trip accuracy. The detailed definitions for three metrics are provided in Appendix C.
@@ 3.2 Performance Comparison
Top-$k$ Round-Trip Accuracy. To assess the accuracy of our predicted synthesis plans, we utilize the Molecule Transformer [21] as the benchmark reaction prediction model and calculate the top-$k$ round-trip accuracy. We take RetroPrime, Retroformer and Graph2SMILES as our strongv SMILES-based baselines. We also use take graph-based method GraphRetro into comparison. The results are presented in Table 5. The data clearly indicate that our model outperforms all SMILES-based baselines by a considerable extent and even exceeds the well-established graph-based method, GraphRetro. This underscores the efficacy of our unsupervised SMILES alignment mechanism, which enables the model to efficiently leverage substructures from product molecules to construct reactants. This mechanism allows the model to focus more intently on learning reaction mechanisms, thereby yielding more plausible predictive outcomes. In summary, our model has exhibited a robust capacity for generating coherent and efficacious synthesis pathways, specifically tailored for advanced downstream applications such as multi-step retrosynthesis planning.
@@ 3.4 Case Study (Visualization of cross-attention mechanism in transformer with UAling)
To further investigate the impact of the proposed SMILES alignment mechanism on model training, we visualize the cross-attention coefficients of different Transformer decoder layers. The visualization is provided in the supplementary information in Fig. 4 of appendix. From Fig. 4, we can observe significant variations in the cross-attention across different layers. Additionally, the establishment of correspondence does not occur exclusively at certain layers, such as the first or last layer. This suggests that directly imposing supervised signals on the cross attention coefficient [32, 23] for SMILES alignment is not a wise approach, whether applied to all layers or only the last layer. It also verifies our statement in Sec. 2.3 that using unsupervised methods for SMILES alignment does not affect the diversity of attention maps and further has a negative impact on model training and performance. This is why our unsupervised SMILES alignment mechanism achieves better results than supervised SMILES alignment.
@@ 5 Conclusion
We provide the details about our metrics, data preprocess and more visualization results in Appendix as our supplementary infomation.
[…]
Acknowledgements
Acknowledgements are not compulsory. Where included they should be brief. Grant or contribution numbers may be acknowledged.
Please refer to Journal-level guidance for any specific requirements.
@@ 5.8 Consent for publication
Appendix A Notations
Table 7 lists the notations used to facilitate reading.
Table 7: Notations for facilitating reading.
Notation Description
$G=(V,E)$ a graph with nodes $V$ and edges $E$
$rank(a,O)$ the position of atom $a$ in DFS order $O$
$root(G,O)$ the atom (node) of G with the minal rank in $O$
$Smiles(G,O)$ The SMILES of molecule graph $G$ generated via DFS order $O$
$M=(V_{M},E_{M})$ a molecule graphs
$M^{\prime}=(V_{M^{\prime}},E_{M^{\prime}})$ a molecule graph
$\tilde{M}=(V_{M}\cup V_{M^{\prime}},E_{M}\cup E_{M^{\prime}})$ a molecule graph consist of two isolated molecules $M$ and $M^{\prime}$
$O_{M}$ a DFS order of molecule graph $M$
$O_{\tilde{M}}$ a DFS order for molecule graph $\tilde{M}$
$\mathcal{R}$ the set of all reactant molecule graphs
$P=(V_{P},E_{P})$ the product molecule graph
$R=(V_{R},E_{R})$ a single reactant molecule graph
$\mathcal{D}(V_{P}),\mathcal{D}(V_{R})$ the set of all possible DFS order for product molecule graph / a single reactant molecule graph
$O_{P},O_{R}$ a DFS order of product, a DFS order of reactant molecule graphs
$CO(R,O_{P})$ the DFS order of reactant $R$ that has a nearly consistent atomic appearance sequence with $O_{P}$
$OPSmiles(\mathcal{R},O_{P})$ the order-preserving reactants SMILES of $\mathcal{R}$ based on DFS order $O_{P}$
$h^{(k)}_{u}$ the node feature of node $u$ at $k$-th iteration of message passing.
$e^{(k)}_{u,v}$ the edge feature of edge between node $u$ and $v$ at $k$-th iteration of message passing.
$H$ the output of the encoder, i.e. the encoded node features
$Z$ the input token embedding for Transformer decoder
$\hat{H}$ the order-aware node features
$\hat{Z}$ the output token embedding of Transformer decoder
$\hat{T}$ the predicted token logits list
$\hat{t}_{i}$ the $i$-th token logits in the predicted token list
$T$ the ground truth token list
$t_{i}$ the $i$-th ground truth token
Appendix B Dataset
USPTO-50K. The USPTO-50K is of high quality, we only remove the reactions whose products only contain a single atom from the dataset.
USPTO-FULL. The raw data of USPTO-FULL provided by Dai et al has error annotations and other problems. Thus we perform the following steps to clean up the data.
  • remove the reactions where different atoms share the same atom map number.
  • remove the reactions whose product is consist of a single atom.
  • remove the reactions with invalid SMILES. The empty SMILES is considered as invalid SMILES too.
  • remove the reactions where product contains atoms that do not appear in reactants.
  • remove the reactants that all atoms do not appear in the product.
USPTO-MIT. The raw data of USPTO-MIT provided by [11]. The original dataset contains reactions that have multiple product molecules. And the reagents are put together with reactions. Thus we perform the following steps to make the dataset more suitable for retrosynthesis prediction task.
  • remove the reactions whose product is consist of a single atom
  • remove the reagents from the reactants. We consider the molecules whose atoms have no intersection with the product as reagents.
  • remove the reactions with multiple product molecules.
The statistical information of the datasets used in this work is summarized in Table 8.
Table 8: Summary of datasets used in this paper. #Train/#Valid/#Test denotes the number of samples in the training/validation/test set, respectively. #Total is the sum of #Train, #Valid and #Test.
Dataset #Train #Valid #Test #Total
USPTO-50K 40,006 5,001 5,007 50,016
USPTO-MIT 395,498 29,076 38,648 453,222
USPTO-FULL 768,679 96,076 96,015 960,770
Appendix C Evaluation Metrics
Top-$k$ Accuracy. We use the conventional top-$k$ accuracy to evaluate the performance of model. A prediction result is considered as correct if and only if all the reactants are correctly predicted.
Top-$k$ SMILES Validity. As SMILES is considered as correct when it can be identified by RDKit [14]. The top-$k$ SMILES validity is calculated as $\frac{1}{N\times k}\sum_{i=1}^{N}\sum_{j=1}^{k}\mathbbm{1}(\mathrm{SMILES\ is\ valid})$, where $N$ is the number of evaluated samples.
Top-$k$ round-trip Accuracy. There may be multiple methods to synthesize a given product. Evaluating the model’s performance solely based on top-k accuracy may lead to biased results. Therefore, in addition to top-k accuracy, we also employ top-k round-trip accuracy as an extra metric to evaluate the model’s performance. Top-k round-trip accuracy measures the percentage of predicted reactants that can undergo a reaction and yield the given product. To align with our baseline, we use the Molecule Transformer [21] as the forward reaction prediction model. The calculation of top-k round-trip accuracy follows the same approach as described in [32].
Appendix D Visualization of the Cross-Attention Coefficients
Figure 4: Visualization of cross attention over product atoms and predicted tokens of different Transformer decoder layers. The number on the y-axis is the map number of atoms in the product. The reactants atoms that not appear in product is colored red in the x-axis. $\circ$ represents the end token.
Appendix E Implementation details
E.1 Generation of Order-Preserving Reactant SMILES
Given a traversal order $O_{P}$ of product $P$, there might be multiple DFS orders of a reactant molecule that satisfy the condition to be the $O_{P}$-corresponding order. This is because reactants might contain atoms and substructures that are not present in the final product. Thus, we use Algorithm 1 to generate the corresponding order defined in Eq. 2 given product molecule $P$, reactant molecule $R$ and a DFS order $O_{P}$ of $P$, where $am(i)$ represents the atom map number of atom $i$.
Algorithm 1 Get Corresponding Order
Input: product graph $P=(V_{P},E_{P})$, DFS order of product $O_{P}$, reactant $R=(V_{R},E_{R})$.
function get_rank($O_{P}$: DFS order of $P$, $x$: atom, $R$: reactant, $AMRank$: a dict, $Vis$: a set)
     Add $x$ into $Vis$.
     Initialize $min\_son$ as $\infty$.
     for $y\in\mathrm{GET\_NEIGHBOR(x,R)}$ do
         if $y\notin Vis$ then
              GET_RANK($O_{P}$, y, $R$, $AMRank$, $Vis$).
              $min\_son=\min(min\_son,AMRank[am(y)])$.
         end if
     end for
     if $x\in O_{P}$ then
         $AMRank[am(x)]=rank(x,O_{P})$.
     else
         $AMRank[am(x)]=min\_son$.
     end if
end function
function get_CO($CO$: a list, $x$: atom, $R$: reactant, $AMRank$: a dict, $Vis$: a set)
     Add $x$ into $Vis$.
     Append $x$ to $CO$.
     Initialize $N$ as an empty list.
     for $y\in\mathrm{GET\_NEIGHBOR(x,R)}$ do
         if $y\notin Vis$ then
              Append $y$ to $N$.
         end if
     end for
     Sort $N$ in ascending order according to $AMRank[am(x)],x\in N$.
     for $y\in N$ do
         GET_CO($CO$, $y$, $R$, $AMRank$, $Vis$).
     end for
end function
Initialize $Vis$ as an empty set.
Initialize $CO$ as an empty list.
Initialize $AMRank$ as a empty dict.
Set $root$ as the atom $x\in V_{R}$ that $am(x)=am(root(P,O_{P}))$.
GET_RANK($O_{P}$, $root$, $R$, $AMRank$, $Vis$)
Reset $Vis$ as an empty set.
GET_CO($CO$, $root$, $R$, $AMRank$, $Vis$).
Output: $CO(R,O_{P})=CO$
E.2 Model Implementation details (Training and Inference)
We implement our model based on Pytorch 1.13 [18] and torch_geometric 2.2.0 [7]. For model for USPTO-50K, we set the hidden size as 512, encoder layer as 8, decoder layer as 8 and the number of attention heads as 8. The dropout ratio is set as 0.1. For model for USPTO-FULL, we set the hidden size as 768, encoder layer as 8, decoder layer as 8 and the number of attention heads as 12. The dropout ratio is also set as 0.1. The highest learning rate of each model as set as 1.25e-4. We slowly increase our learning rate to the highest in the first few epochs and slowly decrease it using exponential decay. The model is trained with Adam optimizer [13].
It is important to note that a prescribed DFS order may implicitly capture the distribution of reactant sub-groups within the product, which could inadvertently result in information leakage. To mitigate this issue, we employ a DFS order during the inference process that generates a canonical SMILES representation for the product molecule, ensuring a standardized input that precludes information leakage.
Appendix F Initial Node Features and Edge Features
We use the atom and bond descriptors provided by Open Graph Benchmark [9]. There are nine atom descriptors including atomic number, chirality, formal charge and other properties that can be calculated by RDKit [14]. And three bond descriptors are used, containing bond type, bond stereochemistry as well as whether the bond is conjugated. All the descriptors provide integer outputs. Thus we set up learnable embedding tables for each descriptor. The initial node features and edge features are then obtained by summing up the embedding corresponding to the output of each descriptor.

### v2

@@ Abstract
Motivation Retrosynthesis planning poses a formidable challenge in the organic chemical industry, particularly in pharmaceuticals. Single-step retrosynthesis prediction, a crucial step in the planning process, has witnessed a surge in interest in recent years due to advancements in AI for science. Various deep learning-based methods have been proposed for this task in recent years, incorporating diverse levels of additional chemical knowledge dependency.
Results This paper introduces UAlign, a template-free graph-to-sequence pipeline for retrosynthesis prediction. By combining graph neural networks and Transformers, our method can more effectively leverage the inherent graph structure of molecules. Based on the fact that the majority of molecule structures remain unchanged during a chemical reaction, we propose a simple yet effective SMILES alignment technique to facilitate the reuse of unchanged structures for reactant generation. Extensive experiments show that our method substantially outperforms state-of-the-art template-free and semi-template-based approaches. Importantly, our template-free method achieves effectiveness comparable to, or even surpasses, established powerful template-based methods.
Scientific contribution We present a novel graph-to-sequence template-free retrosynthesis prediction pipeline that overcomes the limitations of Transformer-based methods in molecular representation learning and insufficient utilization of chemical information. We propose an unsupervised learning mechanism for establishing product-atom correspondence with reactant SMILES tokens, achieving even better results than supervised SMILES alignment methods. Extensive experiments demonstrate that UAlign significantly outperforms state-of-the-art template-free methods and rivals or surpasses template-based approaches, with up to 5% (top-5) and 5.4% (top-10) increased accuracy over the strongest baseline.
@@ 1 Introduction
Retrosynthesis prediction is a crucial task in organic chemistry, aiding in finding efficient synthetic pathways from target molecules to accessible starting materials. Despite significant advancements in chemical synthesis technology, it still remains a challenge in industries like pharmaceuticals. The extensive search space and the incomplete understanding of chemical reaction mechanisms make retrosynthesis prediction difficult, even for experienced chemists. To address this issue, computer-assisted synthetic planning (CASP) has gained increasing attention in recent years, starting from the seminal work by Corey. This paper focuses on single-step retrosynthesis prediction, which is the fundamental step in CASP. It aims to predict the reactants that can lead to a given product molecule through a single reaction step.
Various deep-learning-based single-step retrosynthesis prediction methods have been proposed in recent years. These methods can be broadly classified into three groups based on their dependency on additional chemical knowledge: template-based, semi-template-based and template-free methods. Template-based methods [9, 37, 7, 5] require an extra database of reaction templates. They frame the retrosynthesis prediction as a classification or retrieval problem for reaction templates suitable for the given product molecule to be synthesized. Among these solutions, Retrosim [7] utilizes molecular similarity to rank reaction templates; LocalRetro [5] and GLN [9] use graph neural networks to model the relationship between reaction templates and molecules to predict the most suitable reaction template; RetroKNN [37] further improves upon LocalRetro by addressing the issue of data imbalance using K-nearest neighbors (KNN). Template-based methods have strong interpretability and can accurately predict reactants. However, these methods are often unable to cover all cases and suffer from poor scalability due to limitations imposed by the template database.
@@
Semi-template-based methods to a certain extent are more in line with chemical intuition. However, these methods increase the complexity of inference and training as they break down retrosynthesis into two subtasks. Failures in synthon prediction directly affect subsequent reactant completion and overall performance. Besides, methods based on leaving group necessitates an extra leaving group database. This requirement, akin to template-based approaches, imposes limitations on the model’s scalability.
@@ 2.2 EGAT+
The initial node features $h_{u}^{(0)}$ and edge features $e_{u,v}^{(0)}$ are determined via several chemical property descriptors, whose details are shown in Supplementary Sec. 6. After $K$ iterations of message passing, we can obtain the encoded features $h^{(K)}_{u}$ of all atoms and make up the output $H\in\mathbb{R}^{V_{P}\times d}$ of the encoder, where $d$ denotes the embedding size.
@@ 2.3 SMILES Alignment
For further discussion, we denote the order-preserving reactant SMILES given the reactant molecules $\mathcal{R}$ and a DFS order $O$ of product as $OPSmiles(\mathcal{R},O)$. An example of the process to generate order-preserving reactants SMILES is shown in Fig. 2. The detailed implementations are presented in Supplementary Sec. 5.1.
@@ 3.1 Evaluation Protocol
Benchmark Datasets. We adopt three datasets for evaluation: (1) USPTO-50K consists of 50,016 atom-mapped reactions grouped into 10 different classes; (2) USPTO-FULL comprises 1,013,118 atom-mapped reactions without any reaction class information. (3) USPTO-MIT consists of 479,035 atom-mapped reactions without any reaction class information. To ensure a fair comparison, we adopt the same training/validation/test splits as those in a previous study [9] for USPTO-50K and USPTO-FULL datasets. The training/validation/test splits is aligned with the previous study [14]. The detailed data processing procedure and the statistical information of the processed dataset are presented in Supplementary Sec. 2.
Metrics. We utilize the following three evaluation metrics for evaluation: top-$k$ accuracy, top-$k$ SMILES validity, and top-$k$ round-trip accuracy. The detailed definitions for three metrics are provided in Supplementary Sec. 3.
@@ 3.2 Performance Comparison
Top-$k$ Round-Trip Accuracy. To assess the accuracy of our predicted synthesis plans, we utilize the Molecule Transformer [23] as the benchmark reaction prediction model and calculate the top-$k$ round-trip accuracy. We take RetroPrime, Retroformer and Graph2SMILES as our strongv SMILES-based baselines. We also use take graph-based method GraphRetro into comparison. The results are presented in Table 5. The results clearly indicate that our model outperforms all SMILES-based baselines by a considerable extent and even exceeds the well-established graph-based method, GraphRetro. This underscores the efficacy of our unsupervised SMILES alignment mechanism, which enables the model to efficiently leverage substructures from product molecules to construct reactants. This mechanism allows the model to focus more intently on learning reaction mechanisms, thereby yielding more plausible predictive outcomes. In summary, our model has exhibited a robust capacity for generating coherent and efficacious synthesis pathways, specifically tailored for advanced downstream applications such as multi-step retrosynthesis planning.
@@ 3.4 Case Study (Visualization of cross-attention mechanism in transformer with UAling)
To further investigate the impact of the proposed SMILES alignment mechanism on model training, we visualize the cross-attention coefficients of different Transformer decoder layers. The visualization is provided in Supplementary Fig. 1 of Supplementary Information. From Supplementary Fig. 1, we can observe significant variations in the cross-attention across different layers. Additionally, the establishment of correspondence does not occur exclusively at certain layers, such as the first or last layer. This suggests that directly imposing supervised signals on the cross attention coefficient [34, 25] for SMILES alignment is not a wise approach, whether applied to all layers or only the last layer. It also verifies our statement in Sec. 2.3 that using unsupervised methods for SMILES alignment does not affect the diversity of attention maps and further has a negative impact on model training and performance. This is why our unsupervised SMILES alignment mechanism achieves better results than supervised SMILES alignment.
3.5 Case Study (Multi-step Retrosynthetic Pathway Planning)
Figure 4: Multistep retrosynthesis predictions by our method. (a) Mitapivat (b) Pacritinib citrate (c) Daprodust. The reaction centers and leaving groups are highlighted in different colors. The pathway pf molecule (a) and (b) come from literature, while the last one is verified by chemical experts.
To explore the suitability of our model for multi-step retrosynthetic pathway planning, we select three distinct molecules as targets for synthetic route design, and the synthesis routes are obtained through iterative calls to our UAlign model, which is trained with the USPTO-FULL dataset. The predicted pathways are summarized in Fig. 4.
The first case study involves Mitapivat, a compound approved for the treatment of hereditary hemolytic anemias in February 2022 [1]. Our model successfully predicted the five-step synthetic route reported in [3], with each step consistently ranked within the top-2 predictions. The first step entails an amide coupling reaction, which our model placed at rank 2, yielding the reactants 1-(cyclopropylmethyl)piperazine (compound 2) and 4-(quinoline-8-sulfonamido)benzoic acid (compound 3). Notably, at the initial step, our model also proposed an alternative synthesis method utilizing the Borch Reductive Amination, which was ranked at the first and is consistent with the synthetic route delineated by Saunders et al. Subsequently, for the synthesis of 4-(quinoline-8-sulfonamido)benzoic acid, the model precisely executed a functional group protection strategy during the second step and accurately anticipated the subsequent formation of the sulfonamide, effectively deconstructing the target molecule into readily available precursors. For the synthesis of 1-(cyclopropylmethyl)piperazine, the model strategically protected the amine functional group with a tert-butyloxycarbonyl moiety at the outset and, in the ultimate step, prognosticated the N-alkylation reaction with a top-ranking accuracy. This example illustrates our model’s capability to uncover diverse reaction centers in molecular retrosynthetic design and to generate plausible reactant combinations based on these insights.
The second case under scrutiny is Pacritinib, an orally bioavailable and isoform selective JAK-2 inhibitor for the treatment of patients with myelofibrosis, which received FDA approval on February 28, 2022 [41]. As shown in Figure 4(b), our model successfully delineates a eight-step synthesis, as described in the literature [4], tracing the synthetic pathway from commercially available 5-nitrosalicylaldehyde and 2,6-dichloropyrimidine to the final product. The initial step of the reverse synthesis is olefin metathesis, ranking the first in order of likelihood, followed by another rank-2 aromatic substitution of 4-(3-((allyloxy)methyl)phenyl)-2-chloropyrimidine (compound 12) and 3-((allyloxy)methyl)-4-(2-(pyrrolidin-1-yl)ethoxy)aniline (compound 13). Subsequently, synthesis of 4-(3-((allyloxy)methyl)phenyl)-2-chloropyrimidine was correctly identified via continuous allyl substitution and Suzuki cross-coupling reaction as the top and the second choices. The reverse synthesis of 3-((allyloxy)methyl)-4-(2-(pyrrolidin-1-yl)ethoxy)aniline was reduction of the nitro group, followed by another allyl substitution. In the final step, the model’s highest probability prediction was reduction of the aldehyde group, followed by a nucleophilic substitution. Despite the synthesis route involving a considerable number of steps and encompassing a variety of reaction types, our model successfully and accurately predicted each step within the top-2 choices. This accomplishment signifies the robustness and efficacy of our model in the context of retrosynthetic analysis.
The final case is Daprodustat, the first oral hypoxy-inducing factor prolyl hydroxylase inhibitor (HIF-PHI) for the treatment of renal anemia caused by chronic kidney disease (CKD) [11]. This novel compound received approval for market release from the FDA on the 1st of February, 2023 [41]. Our model predicted the three-step synthetic route. The first step reports the hydrolysis of ester at rank 3, which is aligned with the route provided by Duffy et al. Although next two steps provided by our method do not exist in the literature, there are all explainable. The synthesis of ethyl (1,3-dicyclohexyl-2,4,6-trioxohexahydropyrimidine-5-carbonyl)glycinate (compound 21) was identified via dehydration condensation of 1,3-dicyclohexyl-2,4,6-trioxohexahydropyrimidine-5-carboxylic acid (compound 23) and ethyl glycinate (compound 22) as the top choice, which avoided using toxic ethyl isocyanatoacetate reported in literature. In the final step, the model’s highest probability prediction was amidation of ester, resulting in cost-effective and readily accessible starting materials. This case demonstrates the robust extrapolative capacity of our model, highlighting its potential to generate synthetic routes that surpass those documented in the literature.
@@ 5 Conclusion
We provide the details about our metrics, data preprocess and more visualization results in another pdf file as Supplementary Information.
@@ 5.1 Funding
This work was supported by the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), the SJTU AI for Science platform, and the Fundamental Research Funds for the Central Universities.
@@ 5.5 Author contribution
K.Z. was responsible for the code implementation, algorithm design, and the majority of the manuscript writing. X.Z. conducted the Ablation Study. Y.Z. and B.Y. managed multi-step Retrosynthetic Pathway Planning. F.N. handled all visualizations. The remaining authors contributed to the polishing of the article. Y.X., Y.J. and X.Y. supervised the research.